# Autonomy Architectures for Safe Planning in Unknown Environments Under Budget Constraints

### v1

@@ Abstract
Mission planning can often be formulated as a constrained control problem under multiple path constraints (i.e., safety constraints) and budget constraints (i.e., resource expenditure constraints). In a priori unknown environments, verifying that an offline solution will satisfy the constraints for all time can be difficult, if not impossible. Our contributions are as follows: 1) We propose an online method, building on our previous work gatekeeper, to guarantee safety and satisfy budget constraints of the system trajectory at all times throughout a mission. 2) Next, we prove that our algorithm is recursively feasible and correct. 3) Finally, instead of using a heuristically designed backup controller, we propose a sampling-based method to construct backup trajectories that both minimize resource expenditure and reach budget renewal sets, in which path constraints are satisfied and the constrained resources are renewed. We demonstrate our approach in simulation with a fixed-wing UAV in a GNSS-denied environment with a budget constraint on localization error that can be renewed at visual landmarks. [Code]
@@ I Introduction
Robotic and autonomous systems often operate with limited knowledge of their environment and limited sensing capabilities, which makes designing missions with strict safety requirements challenging. In addition, these systems have a limited budget on resources that are expended throughout a mission. In this paper, we treat safety requirements in the form of a set of states in which the robot must always remain, called the safe set. Specifically, we consider the case where the safe set is not fully known a priori and must be built online with information from the sensor measurements. Online methods require guarantees on recursive feasibility to remain safe if planning methods fail to find a trajectory.
While full trajectory optimization methods exist to handle constrained problems [1], they are often non convex and computationally intractable to solve online. Path planning techniques [2] can generate trajectories quickly based on approximated dynamics, but the solutions may not be dynamically realizable and safety guarantees cannot be guaranteed.
@@
Constraints on finite resources have been studied under various names, including integral constraints [3, 4], budget constraints [5, 6], and cost constraints [7, 8]. Examples include constraints on battery charge [9], localization error [10], or time spent visible to an enemy observer [11]. We consider a generalized case of the budget-constrained planning problem where resources can be renewed at specific regions in the state space.
Various methods to enforce budget constraints have been designed in previous works, ranging from solutions for specific systems and types of missions to general solutions solved offline. An example of a specific application of localization error constrained planning for UAVs is in [8], in which a greedy multi-objective motion planning strategy is used to navigate with little knowledge of the environment, making use of cellular signals to localize the robot.
In [12], a control barrier function (CBF) is used to enforce battery charge constraints, but it assumes a simple, differentiable battery model and is limited to the class of systems with single-integrator dynamics. The methods in [5, 3] solve full optimal control problems with budget constraints and resets in discrete and continuous state spaces, but these approaches rely on large-scale graph searches or solving viscosity equations derived from dynamic programming, which are computationally expensive and better suited for offline planning in known environments.
We note that none of these solutions explicitly handle the combination of general nonlinear dynamics, limited environmental knowledge, and non-trivial budget expenditure dynamics, all while guaranteeing safety. In addition, some of these methods involve mission- and constraint-specific design of the planning modules, which reduces modularity and adaptability of the system to additional types of constraints and missions. To that end, we aim to build a layered control architecture with modular components [13].
Our proposed solution to the problem of ensuring safety and budget constraint guarantees for nonlinear systems in unknown environments builds on our previous work (gatekeeper [14]) to incorporate budget constraints and provides a method for constructing backup trajectories. gatekeeper is a modular framework that ensures safety by recursively constructing a candidate trajectory consisting of a nominal trajectory, followed up to some switching time, and a backup trajectory that guarantees constraint satisfaction thereafter. Recently, gatekeeper has been applied in the eware filter which is designed specifically for battery life constraints, where energy is renewed at a charging station [9, 15]. Our solution further generalizes gatekeeper to enforce a budget constraint on a resource which is expended during the mission and can be renewed in designated regions.
In summary, this paper has three main contributions:
  • We propose an architecture to guarantee safety and satisfy budget constraints while following a nominal trajectory in unknown environments, based on gatekeeper.
  • We prove recursive feasibility and correctness, ensuring that system trajectories satisfy all constraints at all times.
  • We propose a sampling-based method called ReRoot to construct a backup policy that minimizes resource expenditure while ensuring the system reaches budget renewal sets, sets where safety constraints hold and resources are replenished.
@@
The paper is organized as follows. In Section II, we introduce the problem formulation. Section III describes the method overview of gatekeeper and ReRoot. Section IV discusses a simulation case study of a UAV in a GNSS-denied environment navigating to a goal with safety and budget constraints related to visual odometry.
@@ II-A System Modeling
$\begin{cases}\dot{b}=L(x,u),\quad&x\notin\mathcal{R}\\ b^{+}=b_{\textup{reset}},\quad&x\in\mathcal{R}\end{cases}$ (3)
with initial condition
$b(t_{0})=\begin{cases}b_{0},\quad&x\notin\mathcal{R}\\ b_{\textup{reset}},\quad&x\in\mathcal{R},\end{cases}$ (4)
where $L:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ is piecewise continuous and $b^{+}=b_{\textup{reset}}\geq 0$ is the reset value of the budget state after the jump. We assume that $L$ is known, but $\mathcal{R}$ is not assumed to be fully known.
II-B Set Representation of the Environment
We aim to satisfy the following constraints:
$x(t)\in\mathcal{S},$ $\quad\forall t\geq t_{0}$ (5)
$b(t)\leq B,$ $\quad\forall t\geq t_{0}$ (6)
where $\mathcal{S}\subset\mathcal{X}$ is the set of states that satisfy mission safety requirements and $B$ is the budget constraint that must not be exceeded. We assume that $\mathcal{S}$ has a nonempty interior and that $B>b_{\textup{reset}}$.
We assume $\mathcal{S}$ is not fully known, but can be sensed online from sensor measurements. We introduce the known subset of the safe set at time $t_{k}$, $\mathcal{F}_{k}\subset\mathcal{S}$, where $k\in\mathbb{N}$. Similarly, we denote the known renewal set as $\mathcal{R}_{k}\subset\mathcal{R}\subset\mathcal{S}$, where we have assumed the renewal set is safe.
@@ II-B Set Representation of the Environment
$\mathcal{F}_{k}\subset\mathcal{F}_{k+1}\quad\forall k\in\mathbb{N}$ (7)
$\mathcal{R}_{k}\subset\mathcal{R}_{k+1},~{}\forall k\in\mathbb{N}.$ (8)
@@
A set $\mathcal{B}\subset\mathcal{X}$ is a backup set if there exists a controller $\pi:\mathcal{X}\to\mathcal{U}$ such that for the closed-loop system (2), $\forall x(t_{0})$ in a neighborhood $\mathcal{N}\subset\mathcal{X}$, there exists a finite time $0\leq T_{B}(x(t_{0}))<\infty$ such that $\mathcal{B}$ is reachable at $t_{0}+T_{B}$ and controlled invariant thereafter, i.e., $x(t)\in\mathcal{B},~{}\forall t\geq t_{0}+T_{B}$.
@@
We assume that the budget renewal set $\mathcal{R}_{k}$ is a backup set for the closed loop system (2) for all $k\in\mathbb{N}$. Then, from (3), we obtain:
@@ II-C Nominal Trajectory
A trajectory with horizon $T_{H}$ is a piecewise continuous function $p:\mathcal{T}\to\mathcal{X}$ defined on $\mathcal{T}=[t_{0},t_{0}+T_{H}]\subset\mathbb{R}_{\geq 0}$. A trajectory $p$ is dynamically feasible w.r.t. (1) if there exists a piecewise continuous control $u:\mathcal{T}\to\mathcal{U}$ s.t.
[…]
Assumption 2.
We assume there exists a feedback controller $\pi^{T}:\mathcal{X}\times\mathcal{X}\to\mathcal{U}$, where $u(t)=\pi^{T}(x(t),p(t))$, which tracks a given trajectory $p$ by minimizing the error $\left\|x(t)-p(t)\right\|$ for $t\geq t_{0}$. We refer to this policy as the tracking controller.
We assume there exists a nominal planning module that computes a nominal trajectory that achieves the mission objectives. We denote the nominal trajectory from time $t_{k}$ as $p_{k}^{\textup{nom}}:[t_{k},t_{k}+T_{H}]\to\mathcal{X}$. The nominal trajectory is not required to be dynamically feasible w.r.t. (1). Formally, we state the following assumption:
Assumption 3.
For each $k\in\mathbb{N}$, a nominal trajectory $p_{k}^{\textup{nom}}:[t_{k},t_{k}+T_{H}]\to\mathcal{X}$ is known.
@@
Figure 2: Snapshots of gatekeeper with ReRoot trees rooted at budget renewal sets. The trees are grown in the free set $\mathcal{F}_{k}$, with the path from any node to the root being optimal w.r.t. to the budget cost. The robot discovers new budget renewal sets as it uncovers more of the unknown space.
II-D Problem Statement
@@
Given the known safe set $\mathcal{F}_{k}$ and budget renewal set $\mathcal{R}_{k}$ updated at each $k\in\mathbb{N}$, design an architecture such that the closed-loop system (2) tracks a nominal trajectory $p^{\textup{nom}}_{k}$ while guaranteeing all safety and budget constraints are satisfied, i.e., $x(t)\in\mathcal{S}$ and $b(t)\leq B$ for all $t\geq t_{0}$.
@@ III Method Overview
To address Problem 1, we propose an architecture built around gatekeeper, a safety filter, and ReRoot, a backup policy generation algorithm. In particular, we develop a solution that offers two contributions compared to the earlier version of the gatekeeper framework: 1) it accounts for budget constraints, and 2) we provide a constructive method to generate a backup policy that aims to maximize the time spent tracking the desired trajectories. See Figure 2 for a block diagram showing where gatekeeper fits into the general autonomy stack and Table I for a summary of the nomenclature in this paper.
TABLE I: Nomenclature.
Symbol Definition
$t_{k}$ Start time of gatekeeper iteration $k$
$t_{kS}$ Trajectory switching time $t_{kS}=t_{k}+T_{S}$
$t_{kB}$ Backup trajectory time $t_{kB}=t_{kS}+T_{B}$
$\mathcal{X}\subset\mathbb{R}^{n}$ State space
$\mathcal{U}\subset\mathbb{R}^{m}$ Control input space
$\mathcal{S}\subset\mathcal{X}$ Safe set at time
$\mathcal{F}_{k}\subset\mathcal{S}$ Known safe set at time $t_{k}$
$\mathcal{R}_{k}\subset\mathcal{S}$ Budget renewal set at time $t_{k}$
$\mathcal{G}=(\mathcal{V},\mathcal{E})$ ReRoot graph $\mathcal{G}$ with nodes $\mathcal{V}\subset\mathcal{X}$ and edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$.
$p_{k}^{\textup{nom}}$ $k$-th nominal trajectory on $[t_{k},t_{k}+T_{H}]$
$p_{k}^{\textup{bck},T_{S}}$ backup trajectory starting at time $t_{k}+T_{S}$
$p_{k}^{\textup{can},T_{S}}$ candidate trajectory at time $t_{k}$ with switching time $T_{S}$
$p_{k}^{\textup{com}}$ $k$-th committed trajectory on $[t_{k},t_{kB}]$
@@
First, we summarize the key steps of our method. In the following subsections, we provide mathematical definitions of each type of trajectory along with a proof of correctness. gatekeeper is an online safety verification module that recursively filters the nominal planner output by constructing a trajectory that is guaranteed to satisfy all mission constraints [14]. At the $k$-th iteration of our method, where $k\in\mathbb{N}\setminus 0$, the following steps occur:
@@
  • For all $T_{S}\in[0,T_{H}]$, propagate the nominal trajectory $p_{k}^{\textup{nom}}$ (black line in Figure 2) on the interval $[t_{k},t_{k}+T_{S}]$, then generate a backup trajectory $p^{\textup{bck},T_{S}}_{k}$ that reaches $\mathcal{R}_{k}$ with ReRoot, which forms a set of candidate trajectories.
@@
  • For each candidate trajectory, compute the budget state time history $b(t)$ with (3) and verify that constraints (5) and (6) are satisfied, forming a set of valid trajectories. In Figure 2 when $k=2$, candidate trajectories that enter the red unsafe area are deemed invalid.
@@
  • Select the valid trajectory that maximizes $T_{S}$ as the committed trajectory $p_{k}^{\textup{com}}$. In Figure 2, the committed trajectory is composed of a red nominal portion and a green backup portion, which is generated from ReRoot and will be detailed in the following subsections. If there are no valid trajectories, select the $(k-1)$-th committed trajectory $p^{\textup{com}}_{k-1}$. This step guarantees recursive feasibility.
@@
Our goal is to formally prove that all committed trajectories are guaranteed to satisfy all constraints for all time. To do so, we need to first introduce the definitions below.
III-A Guaranteed Constraint Satisfaction
We formally introduce the notion of a backup trajectory.
@@ III-A Guaranteed Constraint Satisfaction
A trajectory $p:\mathcal{T}\to\mathcal{X}$ over the interval $\mathcal{T}=[t_{k}+T_{S},t_{k}+T_{S}+T_{B}]$ is a backup trajectory $p^{\textup{bck},T_{S}}_{k}:\mathcal{T}\to\mathcal{X}$ to a set $\mathcal{B}\subset\mathcal{X}$ over the interval if the resulting control policy $u=\pi^{T}(x,p^{\textup{bck},T_{S}}_{k})$ renders $\mathcal{B}$ a backup set by Definition 2 at time $t_{kB}=t_{k}+T_{S}+T_{B}$.
Note that the backup trajectory is not required to satisfy constraints (5) and (6) – the validation step of the candidate trajectories will filter out any unsafe backup trajectories, which will be detailed below. As gatekeeper maximizes the time spent tracking the nominal trajectory, the construction of the backup trajectory is key to avoid early switching. In the next subsection, we detail our method for constructing backup trajectories with ReRoot.
The candidate trajectories are constructed from the nominal trajectory and some backup trajectories.
@@
Given a switching time $T_{S}\in[0,T_{H}]$, the solution to the initial value problem
$\dot{p}$ $=f(p,u(t)),~{}p(t_{k})=x_{k},$ (11)
$u(t)$ =$=\begin{cases}\pi^{T}(p(t),p_{k}^{\textup{nom}}(t)),&\quad t\in[t_{k},t_{k}+T_{S})\\ \pi^{T}(p(t),p_{k}^{\textup{bck},T_{S}}(t)),&\quad t\geq t_{k}+T_{S}\end{cases}$ (12)
is the candidate trajectory $p_{k}^{\textup{can},T_{S}}$ defined on $[t_{k},\infty)$.
@@
At the $k$-th iteration, define the set of valid candidate trajectories parameterized by $T_{S}$ as
@@
Finally, we can state the following theorem that gatekeeper guarantees constraint satisfaction for all time.
@@
Suppose $p_{0}^{\textup{can},T_{S}}$ is a dynamically feasible candidate trajectory defined on $[t_{0},\infty)$ that is valid by Definition 6 for some $T_{S}\geq 0$. If $p_{k}^{\textup{com}}$ is determined via Definition 7, then
$p_{k}^{\textup{com}}(t)\in\mathcal{S},\quad\forall t\in[t_{k},\infty),$ (19)
$b(t)\leq B,\quad\forall t\in[t_{k},\infty).$ (20)
Furthermore, if $x(t_{0})=p_{0}^{\textup{com}}(t_{0})$, and the system (1) is driven by $u(t)=\pi^{T}_{k}(x(t),p_{k}^{\textup{com}}(t))$, for $t\in[t_{k},t_{k+1})$, then the closed-loop dynamics (2) satisfy $x(t)\in\mathcal{S}$ and $b(t)\leq B$ for all $t\geq t_{0}$.
@@
The proof is by induction and is based on Theorem 1 in [14].
@@
where $t_{0,B}=t_{0}+T_{S}+T_{B}$. We used the fact that $\mathcal{R}_{0}$ is a backup set by Definition 2 and is thus controlled invariant. Hence, $p_{0}^{\textup{com}}$ satisfies all constraints for all time. Moreover, since $p_{0}^{\textup{com}}$ is dynamically feasible, $x(t)=p_{0}^{\textup{com}}(t)\quad\forall t\in[t_{0},t_{1})$. Then for the budget state solution to (3) with $x(t)$ and the corresponding $u(t)$,
$b(t)\begin{cases}\leq B,\quad t\in[t_{0},t_{0B})\\ =b_{\textup{reset}}\leq B,\quad t\geq t_{0B}\end{cases}$
$\iff b(t)\leq B,\quad\forall t\geq t_{0}.$
@@
Case 2: If $\mathcal{I}_{k+1}=\emptyset$, the committed trajectory is unchanged

### v2

@@ Abstract
Mission planning can often be formulated as a constrained control problem under multiple path constraints (i.e., safety constraints) and budget constraints (i.e., resource expenditure constraints). In a priori unknown environments, verifying that an offline solution will satisfy the constraints for all time can be difficult, if not impossible. We present ReRoot, a novel sampling-based framework that enforces safety and budget constraints for nonlinear systems in unknown environments. The main idea is that ReRoot grows multiple reverse RRT* trees online, starting from renewal sets, i.e., sets where the budget constraints are renewed. The dynamically feasible backup trajectories guarantee safety and reduce resource expenditure, which provides a principled backup policy when integrated into the gatekeeper safety verification architecture. We demonstrate our approach in simulation with a fixed-wing UAV in a GNSS-denied environment with a budget constraint on localization error that can be renewed at visual landmarks. [Code]
@@ I Introduction
Robotic and autonomous systems often operate with limited knowledge of their environment and limited sensing capabilities, which poses challenges to guaranteeing their safe operation. Safety requirements are typically given in the form of a set of states in which the robot must always remain, called the safe set. In unknown environments, the safe set is not fully known a priori and must be built online with information from the sensor measurements.
@@
In addition to instantaneous safety requirements, many robotic systems have a limited budget on resources or quantities that are depleted throughout a mission. Constraints on finite resources have been studied under various names, including integral constraints [1], budget constraints [2, 3], and cost constraints [4]. Previously studied examples include constraints on battery charge [5], localization error [6], or time spent visible to an enemy observer [7]. We consider a generalized case of the budget-constrained, safety-critical planning problem where resources can be renewed at specific regions in the state space.
Various offline methods to enforce budget constraints have been introduced. The methods in [1, 2] solve trajectory-optimization problems with renewable budget constraints. However, these approaches rely on searching over large-scale graphs or solving partial differential equations derived from dynamic programming, which is computationally expensive and better suited for offline planning in known environments. An offline reinforcement-learning approach is studied in [8], where budget constraints for high-dimensional, nonlinear systems are incorporated in the training process, however no strict guarantees of constraint satisfaction are obtained.
In this paper, we instead focus on the problem of online safety and budget constraint satisfaction, where lack of knowledge of the environment requires constant replanning and infinite horizon guarantees. Previous work on online budget constraint satisfaction used simplified robot and budget dynamics [9, 4]. The solutions were designed for specific use cases, limiting their general applicability.
Our approach relies on a run-time safety filter. Many safety-critical robotic systems apply these filters with a backup or recovery policy that reaches an invariant backup set to guarantee safety over an infinite horizon [10]. For systems with instantaneous safety constraints, finding a backup set is often trivial, as most robots can stop in place or return to a known safe configuration. Typical backup policies include braking maneuvers, model predictive control approaches, and learned RL policies [11, 12, 13, 14]. However, to guarantee budget constraint satisfaction, the design of backup policies is more challenging because the budget must be renewed in the backup set, which may only occur in limited subsets of the state space. Additionally, the trajectories to reach the backup sets must satisfy all safety constraints, which forms the backup trajectory generation problem.
To the best of the authors’ knowledge, there are no solutions that explicitly guarantee safety and budget constraint satisfaction with limited environmental knowledge and general, non-trivial budget expenditure dynamics. This paper aims to narrow this gap by proposing a recursively feasible algorithm for the online trajectory generation of nonlinear systems under budget and safety constraints in unknown environments. The key idea and contribution is the sampling-based method called ReRoot, which efficiently generates backup trajectories to minimize resource expenditure while ensuring that the system reaches budget-renewal sets, where resources are replenished. The algorithm is demonstrated on a UAV that uses visual odometry to reach a goal in an unknown, GNSS-denied environment, with guaranteed satisfaction of budget and safety constraints.
@@
The paper is organized as follows. In Section II, we introduce the problem formulation. Section III describes the proposed method, highlighting the safe trajectory planner and ReRoot. Section IV discusses a simulation case study of a UAV in a GNSS-denied environment navigating to a goal with safety and budget constraints related to visual odometry.
@@ II Problem Formulation
[…]
where $L:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ is piecewise continuous and $b^{+}=b_{\textup{reset}}\geq 0$ is the reset value of the budget state after the jump. We assume that $L$ is known, but $\mathcal{R}$ is not fully known. We aim to satisfy the following constraints:
$x(t)\in\mathcal{S}\land b(t)\leq B,\quad\forall t\geq t_{0},$ (4)
where $\mathcal{S}\subset\mathcal{X}$ is the set of states that satisfy safety requirements and $B$ is the budget constraint that must not be exceeded. We assume that $B>b_{\textup{reset}}$, that $\mathcal{S}$ has a nonempty interior, and that it is not fully known, but can be sensed online from sensor measurements. We introduce the known subset of the safe set at time $t_{k}$, $\mathcal{F}_{k}\subset\mathcal{S}$, where $k\in\mathbb{N}$. Similarly, we denote the known renewal set as $\mathcal{R}_{k}\subset\mathcal{R}\subset\mathcal{S}$, where we have assumed the renewal set is safe.
@@
$\mathcal{F}_{k}\subset\mathcal{F}_{k+1},~\mathcal{R}_{k}\subset\mathcal{R}_{k+1},\quad\forall k\in\mathbb{N}.$ (5)
@@
A set $\mathcal{B}\subset\mathcal{X}$ is a backup set if there exists a controller $\pi^{B}:\mathcal{X}\to\mathcal{U}$ such that for the closed-loop system (2), $\forall~x(t_{0})$ in a neighborhood $\mathcal{N}\subset\mathcal{X}$, there exists a finite time $0\leq T_{B}(x(t_{0}))<\infty$ such that $\mathcal{B}$ is reachable at $t_{0}+T_{B}$ and controlled invariant thereafter, i.e., $x(t)\in\mathcal{B},~\forall t\geq t_{0}+T_{B}$.
@@
We assume that the budget renewal set $\mathcal{R}_{k}$ is a backup set for the closed loop system (2) for all $k\in\mathbb{N}$. Then, from (3), the budget constraint is satisfied for all time in $\mathcal{R}_{k}$:
@@
A trajectory with horizon $T_{H}$ is a pair of functions $p:\mathcal{T}\to\mathcal{X}$ and $u:\mathcal{T}\to\mathcal{U}$ defined on $\mathcal{T}=[t_{0},t_{0}+T_{H}]\subset\mathbb{R}_{\geq 0}$ that satisfy
$\dot{p}(t)=f(p(t),u(t))\quad\forall t\in\mathcal{T}.$ (7)
The set of all trajectories at $t\in\mathbb{R}$ starting from $x\in\mathcal{X}$ is
$\Phi(t,x)=\{(p,u)~|~p(t)=x\land(p,u)\text{ is a trajectory}\}.$ (8)
We assume there exists a nominal planner to generate trajectories that aim to fulfill the mission objectives.
Definition 4 (Nominal Trajectory).
At time $t_{k}\in\mathbb{R}$ at state $x_{k}\in\mathcal{X}$, the planner generates the nominal trajectory $(p_{k}^{\textup{nom}},u_{k}^{\rm nom})\in\Phi(t_{k},x_{k})$ defined over the interval $\mathcal{T}=[t_{k},t_{k}+T_{H}]$.
@@
Figure 2: Snapshots of gatekeeper with ReRoot trees rooted at budget renewal sets. The trees are grown in the free set $\mathcal{F}_{k}$. The robot discovers new budget renewal sets as it uncovers more of the unknown space.
@@
Given the known safe set $\mathcal{F}_{k}$ and budget renewal set $\mathcal{R}_{k}$ updated at each $k\in\mathbb{N}$, design a safe planning algorithm that recursively constructs a trajectory $p(t)$ that satisfies all safety and budget constraints, i.e., $p(t)\in\mathcal{S}$ and $b(t)\leq B$ for all $t\geq t_{0}$.
@@ III Method Overview
To address Problem 1, we propose a framework that recursively generates trajectories that are guaranteed to satisfy the safety and budget constraints over an infinite horizon. The solution has two components: A) gatekeeper, an algorithm to construct and validate safe trajectories using the nominal trajectory and a backup trajectory, and B) ReRoot, a sampling-based algorithm that allows for efficient online generation of backup trajectories. See Figure 2 for a block diagram showing where the gatekeeper and ReRoot components fit into the general autonomy stack.
III-A Guaranteed Constraint Satisfaction
@@
Here, we list the key steps of our method, and then we provide mathematical definitions of each type of trajectory along with a proof of correctness. We use the gatekeeper framework [16, 17], which recursively filters the output of the nominal planner by constructing a trajectory that is guaranteed to remain in the safe set for all time. We make use of the reformulated budget dynamics in (3) in the gatekeeper framework to augment the definition of the safe set to include the budget constraint. Algorithm 1 details the $k$-th iteration, where $k\in\mathbb{N}\setminus 0$.
1
[…]
3 while $T_{S}\geq 0$ do
4  $p_{k}^{\textup{can},T_{S}},b^{\textup{nom}}\leftarrow\textnormal{{propagate}}(p^{\textup{nom}}_{k},x(t_{k}),[t_{k},t_{k}+T_{S}])$
5
6 $p^{\textup{bck,T_S}}_{k},b^{bck}\leftarrow\textnormal{{GetBackupFromReRoot}}(\mathcal{G}_{k},p_{k}^{\textup{can},T_{S}}(t_{k}+T_{S}),\mathcal{F}_{k})$
7
8 $p^{\textup{can},T_{S}}_{k}\leftarrow\textnormal{{append}}(p^{\textup{can},T_{S}}_{k},p^{\textup{bck,T_S}}_{k})$
9  $b^{-}\leftarrow b^{\textup{nom}}+b^{\textup{bck}}$
10  if $p_{k}^{\textup{can},T_{S}}$ is valid by Definition 7 then
11     $p_{k}^{\textup{com}}=p_{k}^{\textup{can},T_{S}}$
12     return $p^{\textup{com}}_{k}$
13
14 $T_{S}\leftarrow T_{S}-\Delta T$
15
[…]
17 return $p^{\textup{com}}_{k}$
Algorithm 1 gatekeeper with ReRoot
We also summarize the steps below:
@@ 14 $T_{S}\leftarrow T_{S}-\Delta T$
  • For all $T_{S}\in[0,T_{H}]$, propagate the nominal trajectory $p_{k}^{\textup{nom}}$ (black line in Figure 2) on the interval $t_{k},t_{k}+T_{S}]$, then generate a backup trajectory $p^{\textup{bck},T_{S}}_{k}$ that reaches $\mathcal{R}_{k}$ with ReRoot, which forms a set of candidate trajectories.
@@
  • For each candidate trajectory, compute the budget state time history $b(t)$ with (3) and verify that constraints (4) are satisfied, forming a set of valid trajectories. In Figure 2 when $k=2$, candidate trajectories that enter the red unsafe area are deemed invalid.
@@
  • Select the valid trajectory that maximizes $T_{S}$ as the committed trajectory $p_{k}^{\textup{com}}$. In Figure 2, the committed trajectory is composed of a red nominal portion and a green backup portion, which is generated from ReRoot and will be detailed in the following subsections. If there are no valid trajectories, select the (k1)(k-1)-th committed trajectory $p^{\textup{com}}_{k-1}$. This step guarantees recursive feasibility.
@@
To formally prove that all committed trajectories are guaranteed to satisfy all constraints for all time, we need to introduce the notion of a backup trajectory.
@@
A trajectory $(p^{\textup{bck},T_{S}}_{k},u^{\textup{bck},T_{S}}_{k})\in\Phi(t_{k}+T_{S},x(t_{k}+T_{S}))$ over the interval $\mathcal{T}=[t_{k}+T_{S},\infty)$ is a backup trajectory to a set $\mathcal{B}\subset\mathcal{X}$ if the following two conditions hold:
$p^{\textup{bck},T_{S}}_{k}(t_{k}+T_{S}+T_{B})\in\mathcal{B},$ (9)
$u^{\textup{bck},T_{S}}_{k}(t)=\pi^{B}(p^{\textup{bck},T_{S}}_{k}(t))\quad\forall t\geq t_{k}+T_{S}+T_{B}.$ (10)
The backup trajectory is dynamically feasible and enters a backup set within $T_{B}$ seconds after the switching time $T_{S}$. After entering $\mathcal{B}$, the trajectory remains in $\mathcal{B}$ for all time because the control policy $\pi^{B}$ renders $\mathcal{B}$ controlled-invariant.
Note that the backup trajectory is not required to satisfy constraints (4) – the validation step of the candidate trajectories will filter out unsafe backup trajectories. In general, finding a backup trajectory from any initial condition to a backup set is a key step in the gatekeeper framework, which is covered in the next subsection on ReRoot.
A candidate trajectory follows the nominal trajectory then switches to a backup trajectory.
@@
At time $t_{k}\in\mathbb{R}$ and starting from $x_{k}\in\mathcal{X}$, let the nominal trajectory be $(p_{k}^{\textup{nom}},u_{k}^{\rm nom})\in\Phi(t_{k},x_{k})$. Let $t_{kS}=t_{k}+T_{S}$ and $\mathcal{T}_{S}=[t_{k},t_{kS})$. For any switching time $T_{S}\in[0,T_{H}]$, the backup trajectory is $(p_{k}^{\textup{nom}},u_{k}^{\rm nom})\in\Phi(t_{kS},p_{k}^{\rm nom}(t_{kS}))$. A candidate trajectory $(p_{k}^{\textup{can}},u_{k}^{\rm can})\in\Phi(t_{k},x_{k})$ with switching time $T_{S}$ is defined as
$(p_{k}^{\textup{can}}(t),u_{k}^{\rm can}(t))=\begin{cases}(p_{k}^{\textup{nom}}(t),u_{k}^{\rm nom}(t))~&\text{if}~t\in\mathcal{T}_{S},\\ (p_{k}^{\textup{bck},T_{S}}(t),u_{k}^{\textup{bck},T_{S}}(t))~&\text{if}~t\geq t_{kS}.\end{cases}$ (11)
@@
At iteration $k$, the set of valid candidate trajectories parameterized by $T_{S}$ is
@@
The following assumption is required for the inductive recursive feasibility property of gatekeeper.
Assumption 2.
We are given an initial dynamically feasible valid candidate trajectory $p_{0}^{\textup{can},T_{S}}$ which by definition satisfies all safety and budget constraints. [^1] This assumption is not restrictive to satisfy in practice. The initial candidate trajectory can be arbitrarily short and planned within the known safe set, as long as it ends in a backup set. For example, in our case study in Section IV, the initial candidate trajectory is a short out-and-back loop that stays within the initial field of view of the UAV and returns to the controlled-invariant orbit.
Now we can state the following theorem which shows that gatekeeper guarantees constraint satisfaction for all time.
@@
Suppose $p_{0}^{\textup{can},T_{S}}$ is a dynamically feasible candidate trajectory defined on $t_{0},\infty)$ that is valid by Def. 7 for some $T_{S}\geq 0$. If $p_{k}^{\textup{com}}$ is determined via Def. 8, then
$p_{k}^{\textup{com}}(t)\in\mathcal{S}\land b(t)\leq B\quad\forall t\in[t_{k},\infty).$
@@
The proof by induction is based on Theorem 1 in [16].
@@
where $t_{0,B}=t_{0}+T_{S}+T_{B}$. We used the fact that $\mathcal{R}_{0}$ is a backup set by Definition 2 and is thus controlled invariant. Hence, $p_{0}^{\textup{com}}$ satisfies all constraints for all time. Then for the budget state solution to (3) with $p(t)$ and the corresponding $u(t)$,
$b(t)\begin{cases}\leq B,\quad t\in[t_{0},t_{0B})\\ =b_{\textup{reset}}\leq B,\quad t\geq t_{0B}\end{cases}\iff b(t)\leq B,\;\forall t\geq t_{0}.$
@@
Case 2: If $\mathcal{I}_{k+1}=\emptyset$, the committed trajectory is